OpenStudy (anonymous):

Evaluate: lim x^2-x+6/ x-2 x->2

OpenStudy (anonymous):

8/0 = infinity

OpenStudy (anonymous):

The answer is 5, that's what it says in the book. I don't know how they get that.

OpenStudy (anonymous):

s its infinite

OpenStudy (anonymous):

the book is wrong

OpenStudy (anonymous):

just substitute x=2 in all places of x

OpenStudy (anonymous):

I did that and get 0/0

OpenStudy (anonymous):

So it's infinite?

OpenStudy (anonymous):

4-2+6 = 8

OpenStudy (amistre64):

factor the top to see if it cancels the bottom

OpenStudy (anonymous):

Oh.. sorry. It was suppose to be x^2 +

OpenStudy (amistre64):

x^2-x+6 (x+3)(x-2) -------- = ---------- x-2 (x-2)

OpenStudy (amistre64):

as x goes to 2; (x+3) goes to 5

OpenStudy (amistre64):

...... lol, im making up my own problems i think

OpenStudy (anonymous):

x^2-x+6 = (x+3)(x-2) it is not right

OpenStudy (anonymous):

There we go. Thanks Amsitre :)

OpenStudy (anonymous):

Amistre is right.

OpenStudy (amistre64):

there we ALMOST go ;)

OpenStudy (anonymous):

I forgot to factor.

OpenStudy (amistre64):

did you typo the question?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

(x+3)(x-2) = x^2 -x -6

OpenStudy (amistre64):

hah! .... I was right :)

OpenStudy (amistre64):

and that equals x^2 +x-6

OpenStudy (anonymous):

right sorry

OpenStudy (amistre64):

typing and computing are two seperate brain activities ;)

OpenStudy (anonymous):

lol